OpenStudy (jjuden):

. Multiply and simplify: 3xy^4z^2 • 5x^2yx

OpenStudy (anonymous):

combine x terms and y terms if base is same then power will add up in multiplication

OpenStudy (jjuden):

a 8x4y5z2 b 15x2y4z2 c 15x4y5z2 d 15(xyz)11

OpenStudy (anonymous):

so here 3*5*x*x^2*x*y^4*y*z^2=15*x^4*y^5*z^2=15x4y5z2